OpenStudy (anonymous):

What number would you add to the equation below to complete the square? x2 + 18x = 0

OpenStudy (jdoe0001):

do you know what a "perfect square trinomial" is?

OpenStudy (anonymous):

no

OpenStudy (jdoe0001):

\(\bf \large {\color{blue}{ a}}^2\pm2ab+{\color{blue}{ b}}^2\implies ({\color{blue}{ a}}\pm {\color{blue}{ b}})^2\) ^ perfect square trinomial notice the expanded version, specially the MIDDLE term the middle term is 2 * the other 2 guys

OpenStudy (jdoe0001):

so keeping in mind the MIDDLE term, let us group the expression first \(\large{ \bf x^2 + 18x = 0\implies (x^2 + 18x) = 0\\ \quad \\ \implies (x^2 + 18x+{\color{red}{ \square }}^2) = 0}\) so... what do you think we can use there?

OpenStudy (anonymous):

would it be half of 18?

OpenStudy (anonymous):

so 9

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

hmm yes keep in mind that all we're doing is BORROWING FROM ZERO, 0, so if we ADD \(9^2\) we also have to SUBTRACT \(9^2\) thus \(\bf x^2 + 18x = 0\implies (x^2 + 18x) = 0\\ \quad \\ \implies (x^2 + 18x+{\color{red}{ 9 }}^2) -{\color{red}{ 9 }}^2= 0\implies (x+9)^2-9=0\implies (x+9)^2=9\)

OpenStudy (jdoe0001):

woops, got a bit truncated... ok so \(\bf x^2 + 18x = 0\implies (x^2 + 18x) = 0\\ \quad \\ \implies (x^2 + 18x+{\color{red}{ 9 }}^2) -{\color{red}{ 9 }}^2= 0\\ \quad \\\implies (x+9)^2-9=0\implies (x+9)^2=9\)

OpenStudy (anonymous):

so (x+9)^2=9

OpenStudy (jdoe0001):

so (x+9)^2=9 \(\large \checkmark\)

OpenStudy (anonymous):

is our finale answer

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

welp its not one of my answer choices (x+3)^2=9 (x+9)^2=81 (x+18)^2=81 or this fraction one which i dont believe is right

OpenStudy (jdoe0001):

(x+3)^2=9 <--- (x+9)^2=81 (x+18)^2=81

OpenStudy (jdoe0001):

ohh smokes... that's different..hmm

OpenStudy (jdoe0001):

ohh shoot... dohhh, I missed something my bad

OpenStudy (jdoe0001):

\(\bf x^2 + 18x = 0\implies (x^2 + 18x) = 0\\ \quad \\ \implies (x^2 + 18x+{\color{red}{ 9 }}^2) -{\color{red}{ 9 }}^{\huge 2}= 0\)

OpenStudy (jdoe0001):

see the choice now?

OpenStudy (jdoe0001):

\(\bf x^2 + 18x = 0\implies (x^2 + 18x) = 0\\ \quad \\ \implies (x^2 + 18x+{\color{red}{ 9 }}^2) -{\color{red}{ 9 }}^2= 0\\ \quad \\\implies (x+9)^2-9^2=0\implies (x+9)^2=81\)